# FAST: Feature Aware Similarity Thresholding for Weak Unlearning in Black-Box Generative Models

### v1

@@ Abstract
The heightened emphasis on the regulation of deep generative models, propelled by escalating concerns pertaining to privacy and compliance with regulatory frameworks, underscores the imperative need for precise control mechanisms over these models. This urgency is particularly underscored by instances in which generative models generate outputs that encompass objectionable, offensive, or potentially injurious content. In response, machine unlearning has emerged to selectively forget specific knowledge or remove the influence of undesirable data subsets from pre-trained models. However, modern machine unlearning approaches typically assume access to model parameters and architectural details during unlearning, which is not always feasible. In multitude of downstream tasks, these models function as black-box systems, with inaccessible pre-trained parameters, architectures, and training data. In such scenarios, the possibility of filtering undesired outputs becomes a practical alternative.
@@
The primary goal of this study is twofold: first, to elucidate the relationship between filtering and unlearning processes, and second, to formulate a methodology aimed at mitigating the display of undesirable outputs generated from models characterized as black-box systems. Theoretical analysis in this study demonstrates that, in the context of black-box models, filtering can be seen as a form of weak unlearning. Our proposed Feature Aware Similarity Thresholding(FAST) method effectively suppresses undesired outputs by systematically encoding the representation of unwanted features in the latent space. We employ user-marked positive and negative samples to guide this process, leveraging the latent space’s inherent capacity to capture these undesired representations. During inference, we use this identified representation in the latent space to compute projection similarity metrics with newly sampled latent vectors. Subsequently, we meticulously apply a threshold to exclude undesirable samples from the output. Importantly, our proposed methodology exhibits efficacy even in scenarios where the user feedback exclusively consists of negative samples, devoid of any positive examples. The veracity and utility of our approach are empirically substantiated through a series of experiments conducted within a low data ($<100$ samples) regime, featuring the utilization of DCGAN and StyleGAN2 architectures on the MNIST and CelebA-HQ datasets respectively. The code is available at https://github.com/Subhodip123/weak-unlearning-gan
Index Terms:
@@
Machine Unlearning, Black-Box Generative Models, General Adversarial Networks(GANs)
I Introduction
@@
Deep Generative Models have gained widespread popularity for the generation of authentically realistic synthetic data. These models find utility in numerous domains where access to authentic data is limited, exemplified by their application in fields such as medical imaging [1, 2], remote sensing [3, 4], hyperspectral imagery [5, 6] and others[7, 8, 9] and also in various downstream tasks, including but not limited to classification [10, 11], segmentation [12] etc. However, despite their versatile applicability, the presence of undesirable samples within the dataset and the inherent biases ingrained in these datasets [13] contribute to instances where these models produce outputs that are regarded as objectionable, offensive, or even potentially harmful. Thus, recent regulatory measures related to data privacy and protection, such as the European Union’s General Data Protection Regulation (GDPR) [14] and the California Consumer Privacy Act (CCPA) [15], compel organizations to enforce stringent controls on models that have the potential to generate content deemed harmful or offensive. To address this challenge, our focus lies in exploring methodologies that prevent the display of specific output types characterized by undesirable features. In a parallel vein of inquiry, recent research in the domain of Machine Unlearning within generative models [16, 17] has delved into the concept of forgetting, wherein the primary objective revolves around the modification of generative model parameters to ensure the non-generation of undesirable images. It is worth noting that all of the aforementioned studies operate under the assumption that, during the unlearning phase, accessibility to the learned model’s architecture, parameters, and, at times, even the training dataset is readily available. However, there exist scenarios wherein these models are employed as service providers for diverse downstream tasks, functioning as black-box models. In this more stringent context, the pertinent model parameters, architectures, and the underlying dataset remain entirely obscured from the user’s purview. In such a constrained environment, conventional unlearning procedures or the prospect of retraining the model from its inception becomes an unattainable endeavor due to the complete lack of access to model parameters and architectures. Consequently, the question that necessitates elucidation is as follows:
@@
In a black-box scenario, filtering where the task involves screening out undesired outputs while retaining other outputs, is a viable practical solution. This research addresses the formidable challenge of effectively filtering undesired outputs generated by black-box generative models within the feedback-based unlearning framework. Specifically, the user is presented with a curated collection of generated samples stemming from a pre-trained black-box Generative Adversarial Network (GAN). Within this assortment, the user is tasked with distinguishing a subset as undesirable (negative) and another subset as desirable (positive). This task is subject to constraints, notably a limited sample size, typically not exceeding 100 samples. While the concept of employing a binary classifier for filtering appears straightforward, the crux of the challenge resides in training a classifier for this task. The primary hindrance is the scarcity of available samples, making the task inherently intricate. Moreover, conventional filtering mechanisms, which utilize such classifiers, lack awareness of the underlying representation of the undesired feature. Consequently, the development of filtering mechanisms equipped with the capacity to grasp the representation of undesired features becomes a complex and non-trivial endeavor. Adding to the complexity is the requirement to construct a comprehensive and informative representation of the undesired feature. This representation must be meticulously designed to exclusively encapsulate the distinctive characteristics of the undesired feature, without incorporating any extraneous information related to correlated features. The entanglement between the representations of semantic features further complicates this task. For instance, within the CelebA-HQ [18] dataset, the beard feature exhibits substantial correlations with attributes like mustache and male gender. Therefore, an effective filtering mechanism must be equipped with the ability to discern and harness a representation that singularly encapsulates the beard feature, while prudently avoiding the inclusion of information pertaining to other correlated attributes such as mustache and gender.
@@
In this work, we introduce an innovative family of filtering mechanisms, collectively referred to as ‘Feature Aware Similarity Thresholding (FAST)’ meticulously designed to proficiently detect and eliminate outputs featuring undesired attributes while inherently possessing an awareness of the representation of these undesired features. Our methodology, reliant upon user-provided positive and negative samples, effectively exploits the latent space to characterize the representation of the undesired feature. The fundamental rationale underpinning our approach centers on the latent space’s remarkable capacity to encode intricate and interpretable representations of semantic features. This encoded representation encapsulates the essential traits of the undesired feature, thereby endowing the system with the ability to effectively distinguish it from other attributes. The operational framework of our method entails an assessment of the similarity between newly sampled latent vectors and the previously identified latent feature representation corresponding to the undesired attribute. Leveraging the principle that latent vectors encoding similar features are proximate in the latent space, we gauge the alignment between the latent vectors and the undesired feature, facilitating the precise identification of undesired outputs. Subsequently, a well-considered threshold is applied to this similarity metric, leading to the segregation and exclusion of outputs that manifest the undesired feature.
@@
  • This work represents the first instance where filtering is introduced as a form of weak unlearning to block undesired features within black-box generative models.
@@
  • We assess and validate the performance of our proposed method across various GAN architectures, including DCGAN and StyleGAN2, employing datasets such as MNIST and CelebA-HQ, thereby showcasing its effectiveness and versatility in different scenarios and datasets.
II Related Works
II-A Machine Unlearning
@@
Machine unlearning [19, 20] refers to the process of deliberately forgetting specific acquired knowledge or erasing the impact of particular subsets of training data from a trained model. Naive unlearning methods typically entail the removal of undesirable data from the training dataset, followed by retraining the model from scratch. However, this approach becomes computationally prohibitive when unlearning requests are made iteratively for individual data points. Inspired by concerns surrounding privacy protection, [21] introduced methods for data deletion within statistical query algorithms, coining the term ‘machine unlearning’. Unfortunately, these methods are primarily suitable for structured problems and do not extend to complex machine learning algorithms, such as k-means clustering [22] or random forests [23]. Efficient deletion algorithms were devised for the k-means clustering problem, which introduced effective data deletion criteria applicable to randomized algorithms based on statistical indistinguishability. Building upon this criterion, machine unlearning methods are broadly categorized into two main types: exact unlearning [22, 23] and approximate unlearning [24]. Exact unlearning endeavors to completely eradicate the influence of unwanted data from the trained model, necessitating precise parameter distribution matching between the unlearned and retrained models. In contrast, approximate unlearning methods only partially mitigate data influence, resulting in parameter distributions that closely resemble the retrained model, albeit with minor multiplicative and additive adjustments. To eliminate the influence of unwanted data, [25] proposed technique employs parameter perturbation based on cached gradients, offering computational efficiency while increasing memory usage. Other methods [26, 27] have suggested the use of influence functions for this purpose. However, these approaches are computationally demanding due to the necessity of Hessian inversion techniques and are limited to small convex models. To extend the applicability of influence removal techniques to non-convex models like deep neural networks, a scrubbing mechanism [28] was introduced within a classification framework. However, until recently, it remained unclear how these techniques could be applied to unsupervised models, particularly state-of-the-art generative models. In response, [29] introduced an ’Adapt-then-Unlearn’ mechanism designed to unlearn a GAN in a zero-shot setting where the training dataset is inaccessible. Nevertheless, this method does not readily extend to black-box generative models, presenting a challenge in the context of machine unlearning for black-box models.
II-B Generative Adversarial Network
@@
Generative Adversarial Networks (GANs), a prominent class of generative models primarily used for image synthesis, were first introduced in the seminal work by [30]. The GAN framework consists of two neural networks: a Generator (G) network and a Discriminator (D) network. GANs are trained through an adversarial loss function, which strives to minimize the divergence between the distribution of original data and that of generated (fake) data. The Generator network, once trained, takes random latent vectors sampled from a normal distribution as input and produces synthetic data that closely resembles the original data. This process was theoretically elucidated by [30, 31], who demonstrated that the generator aims to minimize the Jensen-Shannon Divergence between the distribution of fake data and that of the original data. Moreover, [32] expanded upon this idea by showcasing that various divergence metrics, with the Jensen-Shannon Divergence being one of them, can be minimized to facilitate the training of effective generative models. They introduced a broader class of models called f-GANs, with the specific choice of the f-divergence metric influencing the optimization process. To further enhance the capabilities of GANs, [33] proposed an information-theoretic extension known as Info-GAN. This innovation allowed GANs to learn a disentangled latent space, significantly improving the interpretability and editability of the generated data. Addressing challenges related to GAN training stability, [31] introduced the idea of training GANs with Wasserstein Distance metrics, which contributed to more stable and reliable training processes. Furthermore, the field of generative models has witnessed the introduction of numerous specialized GAN variants[34] aimed at synthesizing high-quality images. For instance, [35, 36, 37] proposed a style incorporation technique during model training, achieving state-of-the-art results for generating high-resolution images.
II-C Latent Space Representations and Inversion
@@
The latent representation of Generative Adversarial Networks (GANs) often proves to be valuable for auxiliary supervised downstream tasks. Furthermore, this latent space exhibits the capacity to encode a diverse array of semantic features that find utility in tasks such as image manipulation and editing, among others, as demonstrated by [33]. Numerous studies [35, 36, 37] have underscored the latent space’s remarkable representation capabilities, which encompass a wide range of interpretable semantics. For instance, in the context of StyleGAN2, it was observed that the learned $\mathcal{W}$ space possesses greater disentanglement compared to the random noise space $\mathcal{Z}$. This characteristic enhancement in disentanglement contributes to improved editability and perceptual quality in generated images. Additionally, latent spaces often exhibit commendable clustering properties [38], where data with similar semantic features cluster closely within the latent space. To harness the rich structural and representational attributes of the latent space, it is essential to project data back into this space through a process known as inversion [39]. Inverter networks are employed for this purpose, taking data as input and producing corresponding latent vectors, which enable the GAN model to accurately reconstruct the data. Two primary approaches are typically used for inversion: iterative optimization and inference with an encoder [40, 41]. Recent developments in encoder-based methods [42, 43] have yielded impressive outcomes, not only in terms of reconstruction but also in enhancing expressiveness.
III Preliminaries
@@
Consider a specific training dataset $D$ consisting of $m$ i.i.d samples $\{{x_{i}}\}_{i=1}^{m}$, drawn from a distribution $P_{\mathcal{X}}$ over the data space $\mathcal{X}$. A generative model ($G$) trained on $D$ aims to learn the data distribution $P_{\mathcal{X}}$. The pre-trained generative model is denoted as $G_{\theta^{init}}$, with parameters $\theta^{init}\in\Theta$. Based on the outputs of the generative model, the user designates a portion of the data space as undesired, referred to as $\mathcal{X}_{f}$ Therefore, the entire data space can be expressed as the union of $\mathcal{X}_{r}$ and $\mathcal{X}_{f}$, where $\mathcal{X}=\mathcal{X}_{r}\bigcup\mathcal{X}_{f}$. We denote the distributions over $\mathcal{X}_{f}$ and $\mathcal{X}_{r}$ as $P_{\mathcal{X}_{f}}$ and $P_{\mathcal{X}_{r}}$, respectively. The objective is to create a mechanism in which the generative model does not produce outputs within the domain $\mathcal{X}_{f}$. This implies that the generative model should be trained to generate data samples conforming to the distribution $P_{\mathcal{X}_{r}}$. A naive approach to achieving this is by retraining the entire model from scratch using a dataset $D_{r}$. Here, $D_{r}$ is defined as the intersection of $D$ and $\mathcal{X}_{r}$, or equivalently, $D_{r}=D\setminus D_{f}$, where $D_{f}=D\bigcap\mathcal{X}_{f}$. Consequently, any retraining mechanism ($\mathcal{R_{T}}$) results in a new model with parameters $\theta^{r}$ whose output distribution $P_{\theta^{r}}$ seeks to match the distribution $P_{\mathcal{X}_{r}}$.
III-A Machine Unlearning

### v2

@@ Abstract
The heightened emphasis on the regulation of deep generative models, propelled by escalating concerns pertaining to privacy and compliance with regulatory frameworks, underscores the imperative need for precise control mechanisms over these models. This urgency is particularly underscored by instances in which generative models generate outputs that encompass objectionable, offensive, or potentially injurious content. In response, machine unlearning has emerged to selectively forget specific knowledge or remove the influence of undesirable data subsets from pre-trained models. However, modern machine unlearning approaches typically assume access to model parameters and architectural details during unlearning, which is not always feasible. In multitude of downstream tasks, these models function as black-box systems, with inaccessible pre-trained parameters, architectures, and training data. In such scenarios, the possibility of filtering undesired outputs becomes a practical alternative. Our proposed method Feature Aware Similarity Thresholding (FAST) effectively suppresses undesired outputs by systematically encoding the representation of unwanted features in the latent space. We employ user-marked positive and negative samples to guide this process, leveraging the latent space’s inherent capacity to capture these undesired representations. During inference, we use this identified representation in the latent space to compute projection similarity metrics with newly sampled latent vectors. Subsequently, we meticulously apply a threshold to exclude undesirable samples from the output. Our implementation is available at https://github.com/Subhodip123/weak-unlearning-gan
{IEEEImpStatement}
@@
The primary goal of this study is twofold: first, to elucidate the relationship between filtering and unlearning processes, and second, to formulate a methodology aimed at mitigating the display of undesirable outputs generated from black-box Generative Adversarial Networks (GANs). Theoretical analysis in this study demonstrates that, in the context of black-box models, filtering can be seen as a form of weak unlearning. Importantly, our proposed methodology exhibits efficacy even in scenarios where the user feedback exclusively consists of negative samples, devoid of any positive examples. The veracity and utility of our approach are empirically substantiated through a series of experiments conducted within a low data ($<100$ samples) regime, featuring the utilization of DC-GAN and Style-GAN2 architectures on the MNIST and CelebA-HQ datasets respectively.
{IEEEkeywords}
@@
Black-Box Generative Models, General Adversarial Networks(GANs), Machine Unlearning
1 Introduction
@@
Deep generative models have gained widespread popularity for the generation of authentically realistic synthetic data. These models find utility in numerous domains where access to authentic data is limited, exemplified by their application in fields such as medical imaging [1, 2], remote sensing [3, 4], hyper-spectral imagery [5, 6] and others[7, 8, 9] and also in various downstream tasks, including but not limited to classification [10, 11], segmentation [12] etc. However, despite their versatile applicability, the presence of undesirable samples within the dataset and the inherent biases ingrained in these datasets [13] contribute to instances where these models produce outputs that are regarded as objectionable, offensive, or even potentially harmful. Thus, recent regulatory measures related to data privacy and protection, such as the European Union’s General Data Protection Regulation (GDPR) [14] and the California Consumer Privacy Act (CCPA) [15], compel organizations to enforce stringent controls on models that have the potential to generate content deemed harmful or offensive. To address this challenge, our focus lies in exploring methodologies that prevent the display of specific output types characterized by undesirable features. In a parallel vein of inquiry, recent research in the domain of machine unlearning within generative models [16, 17] has delved into the concept of forgetting, wherein the primary objective revolves around the modification of generative model parameters to ensure the non-generation of undesirable images. It is worth noting that all of the aforementioned studies operate under the assumption that, during the unlearning phase, accessibility to the learned model’s architecture, parameters, and, at times, even the training dataset is readily available. However, there exist scenarios wherein these models are employed as service providers for diverse downstream tasks, functioning as black-box models. In this more stringent context, the pertinent model parameters, architectures, and the underlying dataset remain entirely obscured from the user’s purview. In such a constrained environment, conventional unlearning procedures or the prospect of retraining the model from its inception becomes an unattainable endeavor due to the complete lack of access to model parameters and architectures. Consequently, the question that necessitates elucidation is as follows:
@@
In a black-box scenario, filtering where the task involves screening out undesired outputs while retaining other outputs, is a viable practical solution. This research addresses the formidable challenge of effectively filtering undesired outputs generated by black-box generative models within the feedback-based unlearning framework. Specifically, the user is presented with a curated collection of generated samples stemming from a pre-trained black-box Generative Adversarial Network (GAN). Within this assortment, the user is tasked with distinguishing a subset as undesirable (negative) and another subset as desirable (positive). This task is subject to constraints, notably a limited sample size for the user feedback (typically not exceeding 100 samples). While the concept of employing a binary classifier for filtering appears straightforward, the crux of the challenge resides in training a classifier for this task. The primary hindrance is the scarcity of available user feedback samples, making the task inherently intricate. Moreover, conventional filtering mechanisms, which utilize such classifiers, lack awareness of the underlying representation of the undesired feature. Consequently, the development of filtering mechanisms equipped with the capacity to grasp the representation of undesired features becomes a complex and non-trivial endeavor. Adding to the complexity is the requirement to construct a comprehensive and informative representation of the undesired feature. This representation must be meticulously designed to exclusively encapsulate the distinctive characteristics of the undesired feature, without incorporating any extraneous information related to correlated features. The entanglement between the representations of semantic features further complicates this task. For instance, within the CelebA-HQ [18] dataset, the beard feature exhibits substantial correlations with attributes like mustache and male gender. Therefore, an effective filtering mechanism must be equipped with the ability to discern and harness a representation that singularly encapsulates the beard feature, while prudently avoiding the inclusion of information about other correlated attributes such as mustache and gender.
@@
In this work, we introduce an innovative family of filtering mechanisms, collectively referred to as Feature Aware Similarity Thresholding (FAST) meticulously designed to proficiently detect and eliminate outputs featuring undesired attributes while inherently possessing an awareness of the representation of these undesired features. Our methodology, reliant upon user-provided positive and negative samples, effectively exploits the latent space to characterize the representation of the undesired feature. The fundamental rationale underpinning our approach centers on the latent space’s remarkable capacity to encode intricate and interpretable representations of semantic features. This encoded representation encapsulates the essential traits of the undesired feature, thereby endowing the system with the ability to effectively distinguish it from other attributes. The operational framework of our method entails an assessment of the similarity between newly sampled latent vectors and the previously identified latent feature representation corresponding to the undesired attribute. Leveraging the principle that latent vectors encoding similar features are proximate in the latent space, we gauge the alignment between the latent vectors and the undesired feature, facilitating the precise identification of undesired outputs. Subsequently, a well-considered threshold is applied to this similarity metric, leading to the segregation and exclusion of outputs that manifest the undesired feature.
@@
  • This work represents the first instance where filtering is introduced as a form of weak unlearning to block undesired features within black-box Generative Adversarial Networks (GANs).
@@
  • We assess and validate the performance of our proposed method across various GAN architectures, including DC-GAN and Style-GAN2, employing datasets such as MNIST and CelebA-HQ, thereby showcasing its effectiveness and versatility in different scenarios and datasets.
2 Background
2.1 Machine Unlearning
@@
Machine unlearning [19, 20] refers to the process of deliberately forgetting specific acquired knowledge or erasing the impact of particular subsets of training data from a trained model. Naive unlearning methods typically entail the removal of undesirable data from the training dataset, followed by retraining the model from scratch. However, this approach becomes computationally prohibitive when unlearning requests are made iterative for individual data points. Inspired by concerns surrounding privacy protection, [21] introduced methods for data deletion within statistical query algorithms, coining the term machine unlearning. Unfortunately, these methods are primarily suitable for structured problems and do not extend to complex machine learning algorithms, such as k-means clustering [22] or random forests [23]. Efficient deletion algorithms were devised for the k-means clustering problem, which introduced effective data deletion criteria applicable to randomized algorithms based on statistical indistinguishability. Building upon this criterion, machine unlearning methods are broadly categorized into two main types: exact unlearning [22, 23] and approximate unlearning [24]. Exact unlearning endeavors to completely eradicate the influence of unwanted data from the trained model, necessitating precise parameter distribution matching between the unlearned and retrained models. In contrast, approximate unlearning methods only partially mitigate data influence, resulting in parameter distributions that closely resemble the retrained model, albeit with minor multiplicative and additive adjustments. To eliminate the influence of unwanted data, [25] proposed technique employs parameter perturbation based on cached gradients, offering computational efficiency while increasing memory usage. Other methods [26, 27] have suggested the use of influence functions for this purpose. However, these approaches are computationally demanding due to the necessity of hessian inversion techniques and are limited to small convex models. To extend the applicability of influence removal techniques to non-convex models like deep neural networks, a scrubbing mechanism [28] was introduced within a classification framework. However, until recently, it remained unclear how these techniques could be applied to unsupervised models, particularly state-of-the-art generative models. In response, [29] introduced an Adapt-then-Unlearn mechanism designed to unlearn a GAN in a zero-shot setting where the training dataset is inaccessible. Nevertheless, this method does not readily extend to black-box generative models, presenting a challenge in the context of machine unlearning for black-box models.
2.2 Generative Adversarial Network
@@
Generative Adversarial Networks (GANs), a prominent class of generative models primarily used for image synthesis, were first introduced in the seminal work by [30]. The GAN framework consists of two neural networks: a generator network and a discriminator network. GANs are trained through an adversarial loss function, which strives to minimize the divergence between the distribution of original data and that of generated (fake) data. The generator network, once trained, takes random latent vectors sampled from a normal distribution as input and produces synthetic data that closely resembles the original data. This process was theoretically elucidated by [30, 31], who demonstrated that the generator aims to minimize the Jensen-Shannon divergence between the distribution of fake data and that of the original data. Moreover, [32] expanded upon this idea by showcasing that various divergence metrics, with the Jensen-Shannon divergence being one of them, can be minimized to facilitate the training of effective generative models. They introduced a broader class of models called f-GANs, with the specific choice of the f-divergence metric influencing the optimization process. To further enhance the capabilities of GANs, [33] proposed an information-theoretic extension known as Info-GAN. This innovation allowed GANs to learn a disentangled latent space, significantly improving the interpretability and editability of the generated data. Addressing challenges related to GAN training stability, [31] introduced the idea of training GANs with Wasserstein distance metrics, which contributed to more stable and reliable training processes. Furthermore, the field of generative models has witnessed the introduction of numerous specialized GAN variants[34] aimed at synthesizing high-quality images. For instance, [35, 36, 37] proposed a style incorporation technique during model training, achieving state-of-the-art results for generating high-resolution images.
2.3 Latent Space Representations and Inversion
@@
The latent representation of Generative Adversarial Networks (GANs) often proves to be valuable for auxiliary supervised downstream tasks. Furthermore, this latent space exhibits the capacity to encode a diverse array of semantic features that find utility in tasks such as image manipulation and editing, among others, as demonstrated by [33]. Numerous studies [35, 36, 37] have underscored the latent space’s remarkable representation capabilities, which encompass a wide range of interpretable semantics. For instance, in the context of Style-GAN2, it was observed that the learned $\mathcal{W}$ space possesses greater disentanglement compared to the random noise space $\mathcal{Z}$. This characteristic enhancement in disentanglement contributes to improved editability and perceptual quality in generated images. Additionally, latent spaces often exhibit commendable clustering properties [38], where data with similar semantic features cluster closely within the latent space. To harness the rich structural and representational attributes of the latent space, it is essential to project data back into this space through a process known as inversion [39]. Inverter networks are employed for this purpose, taking data as input and producing corresponding latent vectors, which enable the GAN model to accurately reconstruct the data. Two primary approaches are typically used for inversion: iterative optimization and inference with an encoder [40, 41]. Recent developments in encoder-based methods [42, 43] have yielded impressive outcomes, not only in terms of reconstruction but also in enhancing expressiveness.
3 Theory and Analysis
@@
Consider a specific training dataset $D$ consisting of $m$ i.i.d samples $\{{x_{i}}\}_{i=1}^{m}$, drawn from a distribution $P_{\mathcal{X}}$ over the data space $\mathcal{X}$. A generative model ($G$) trained on $D$ aims to learn the data distribution $P_{\mathcal{X}}$. The pre-trained generative model is denoted as $G_{\theta^{init}}$, with parameters $\theta^{init}\in\Theta$. Based on the outputs of the generative model, the user designates a portion of the data space as undesired, referred to as $\mathcal{X}_{f}$. Therefore, the entire data space can be expressed as the union of $\mathcal{X}_{r}$ and $\mathcal{X}_{f}$, where $\mathcal{X}=\mathcal{X}_{r}\bigcup\mathcal{X}_{f}$. We denote the distributions over $\mathcal{X}_{f}$ and $\mathcal{X}_{r}$ as $P_{\mathcal{X}_{f}}$ and $P_{\mathcal{X}_{r}}$, respectively. The objective is to create a mechanism in which the generative model does not produce outputs within the domain $\mathcal{X}_{f}$. This implies that the generative model should be trained to generate data samples conforming to the distribution $P_{\mathcal{X}_{r}}$. A naive approach to achieving this is by retraining the entire model from scratch using a dataset $D_{r}$. Here, $D_{r}$ is defined as the intersection of $D$ and $\mathcal{X}_{r}$, or equivalently, $D_{r}=D\setminus D_{f}$, where $D_{f}=D\bigcap\mathcal{X}_{f}$. Consequently, any retraining mechanism ($\mathcal{R_{T}}$) results in a new model with parameters $\theta^{r}$ whose output distribution $P_{\theta^{r}}$ seeks to match the distribution $P_{\mathcal{X}_{r}}$.
3.1 Machine Unlearning